# Explicitly Disentangled Representations in Object-Centric Learning

### v1

@@ 2.1 Slot Attention
At a given refinement step $t$ of the Slot Attention mechanism, the previous slots $\mathrm{S}^{t-1}$ are mapped as queries and fed to the competitive cross-attention module, which yields one vector $\mathrm{U}_{i}$ for each $\mathrm{Q}_{i}$, referred to as $update$. The updates are then passed as inputs to a shared Gated Recurrent Unit (GRU) Cho et al. (2014), while the associated slot vectors from $\mathrm{S}^{t-1}$ are set as hidden states. The outputs, which are updated hidden states, are transformed by a shared MLP with a residual connection (He et al., 2016), resulting in the new slot vectors $\mathrm{S}^{t}$. The competition between slots for explaining part of the input is induced by an additional normalization in the cross-attention mechanism, performed by using the softmax function over the queries before linearly normalizing over the keys. Formally, the attention coefficients are computed as
@@ 5.2 Disentanglement Analysis
[…]
@@ 5.3 Compositional and Generative Capabilities
Finally, we also explore the compositional and generative capabilities of DISA derived from the disentanglement of texture and shape through simple qualitative experiments. Precisely, as a first investigation, for a given image from the test set we take the representations of the objects, randomly permute their texture parts, and decode the permuted representations. Ideally, the model would output a reconstruction where the shapes, positions, and sizes of the entities are preserved, while their textures are interchanged and adapted to fit the same masks. In the second experiment, we analyze whether this ability extends to the generation of new objects by providing new sampled textures to the objects in a scene. To do so, we average the texture vectors encoded from a given image and sample from a Gaussian distribution centered on that mean vector. For additional results, including those on Multi-dSprites and CLEVRTex, please see Appendix F.3 and F.4.
@@
Figure 3: Compositional capabilities of DISA on (a) Tetrominoes and (b) CLEVR6. The central row contains the input image and the predicted object masks. At the top, the reconstructed image and object textures are shown (empty slots excluded), while the bottom row presents the reconstruction and object textures after interchanging texture vectors between objects.
In Figure 4 we show the generative results of DISA on Tetrominoes and CLEVR6. On Tetrominoes (Figure 3(a)), the decoded textures maintain the original shapes but switch appearance. From the first slot, it is also visible that we can generate coherent out-of-distribution textures, as the dataset only contains objects with plain colors, while the one we obtain mixes red and yellow. Concerning CLEVR6 (Figure 3(b)), we find again that DISA is able to precisely adapt the sampled textures to the fixed shapes, positions, and scales of the objects in the image. Figure 4 shows therefore that the learned representations of DISA also allow for generative capabilities: we can swap in completely new textures for specific objects independently of their shape, location, and scale.
@@
Figure 4: Generative capabilities of DISA on (a) Tetrominoes and (b) CLEVR6. The top row contains the reconstructed input image and object texture. At the bottom, the same visualization is presented after replacing the texture vectors with ones sampled from $\mathcal{N}(\mu_{\text{tex}},0.035)$, with $\mu_{\text{tex}}$ being the mean of the texture vectors encoded from the input image.
@@ Appendix A Sobel Filter
A 2-dimensional convolution operation with $\mathcal{K}_{x}$ and $\mathcal{K}_{y}$ is applied on the input image $\mathcal{X}$, producing two maps of the approximated horizontal and vertical derivatives at each pixel, respectively ${\mathcal{G}_{x}=\mathcal{K}_{x}*\mathcal{X}}$ and ${\mathcal{G}_{y}=\mathcal{K}_{y}*\mathcal{X}}$. Figure 4(b) shows an example of the resulting $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$ of the grayscale image 4(a). To obtain the final filtered input containing the detected edges, the maps are aggregated as
@@ Appendix B Datasets
Tetrominoes consists of $35\times 35$ RGB images presenting 3 “Tetris”-like blocks and a black background. The blocks are characterized by one of 6 possible colors (red, green, blue, yellow, magenta, cyan) and a shape and orientation within 19 possible. Figure 6(a) shows a few examples extracted from the dataset. The tetrominoes can appear anywhere in the image, but cannot overlap with each other. Furthermore, multiple shapes/orientations/colors of the same kind can be included in a single sample. We employ a total of 60K samples for the training set and 320 for the test set, in line with Greff et al. (2019); Locatello et al. (2020).
@@
CLEVR, introduced by Johnson et al. (2017), is a collection of $240\times 320$ images of 3D scenes containing a number of objects ranging from 3 to 10 (excluding the background). There are three types of shapes, i.e. cube, sphere, and cylinder, two possible sizes (small and large), eight colors (gray, red, blue, green, brown, purple, cyan, and yellow), and two materials (shiny “metal” and matte “rubber”). Some samples from CLEVR are presented in Figure 6(c). We also employ a filtered version of this dataset, called CLEVR6, that has a maximum of 6 entities in the scenes, background excluded. The number of training samples, in this case, is 70K, while the test samples are 15K. In all our experiments, a center crop of $192\times 192$ is applied, followed by a resize to a resolution of $128\times 128$.
@@ Appendix E Disentanglement Analysis
(a) Property prediction
(b) Inverse property prediction
Figure 9: Illustration of the two property prediction tasks that we propose to quantitatively study the degree of texture and shape disentanglement in the representations of DISA. (a) Prediction of an object property based on the associated components. For instance, the color from the texture-related latent features. (a) Inverse property prediction task, where properties are predicted based on the “wrong” part of the object representation, e.g., the color from the shape-related components.
@@ E.1 Training and Validation Curves
(e) Material - Training
(f) Material - Validation
@@ Appendix F Additional Experimental Results
F.1 Position and Scale Disentanglement
Figure 13: Qualitative results showing the position and scale disentanglement on DISA. Given an input image (first column), we select two objects and modify their scale/position factors incrementally over 4 steps (second to fifth columns). As visible, texture and shape are consistent across the steps, while position and scale can be independently varied with success. This suggests that the models are able to confine position and scale information in their respective components. Note that the visible limitations are displayed by ISA as well. For example, the missing parts of an object after moving another object covering it. Or even the inconsistent masks produced when incrementing the scale over a certain point. Moreover, it is interesting to see that, on Tetrominoes, scaling an object does not yield completely wrong results. This is not trivial as, for a given shape, only a single variation of its scale is present in the dataset, meaning that the model only sees shapes of fixed size during training.
F.2 Ablation: Impact of Variance Regularization and Sobel Filter on Texture and Shape Disentanglement
@@ F.4 Generative Results
As anticipated, we sample the new textures from a Gaussian distribution centered on the mean of the texture vectors of the objects in a scene, with a standard deviation of 0.035 (manually tuned).
We sample new shapes from a Gaussian distribution centered on the mean of the shape vectors of the foreground objects in a scene, with a standard deviation manually tuned. On Tetrominoes and Multi-dSprites, the standard deviation is set to 0.1, while on CLEVR and CLEVRTex 0.01. Note that this task is quite difficult: first, the number of unique shapes is strongly limited in these synthetic datasets; second, without changing the scale of the reference frame, the generated shape is restricted to one that fits the original frame.
@@
(a) Tetrominoes
(b) Multi-dSprites
(c) CLEVR6
(d) CLEVRTex
Figure 35: Shape generation results. In each of the 4 images, the first pair of rows shows final reconstruction (middle left), object textures (first row), and shapes (second row) obtained after encoding and decoding an input image. The second pair of rows presents final reconstructions (middle left) and individual object textures (first row) and shapes (second row) after sampling new shape vectors.

### v2

@@ 2.1 Slot Attention
At a given refinement step $t$ of the Slot Attention mechanism, the previous slots $\mathrm{S}^{t-1}$ are mapped as queries and fed to the competitive cross-attention module, which yields one vector $\mathrm{U}_{i}$ for each $\mathrm{Q}_{i}$, referred to as $update$. The updates are then passed as inputs to a shared Gated Recurrent Unit (GRU) (Cho et al., 2014), while the associated slot vectors from $\mathrm{S}^{t-1}$ are set as hidden states. The outputs, which are updated hidden states, are transformed by a shared MLP with a residual connection (He et al., 2016), resulting in the new slot vectors $\mathrm{S}^{t}$. The competition between slots for explaining part of the input is induced by an additional normalization in the cross-attention mechanism, performed by using the softmax function over the queries before linearly normalizing over the keys. Formally, the attention coefficients are computed as
@@ 5.2 Disentanglement Analysis
(a) Property prediction
(b) Inverse property prediction
Figure 2: Illustration of the two property prediction tasks that we propose to quantitatively study the degree of texture and shape disentanglement in the representations of DISA. (a) Prediction of an object property based on the associated components. For instance, the color from the texture-related latent features. (a) Inverse property prediction task, where properties are predicted based on the “wrong" part of the object representation, e.g., the color from the shape-related components.
Figure 3: Quantitative results of DISA on the regular and inverse property prediction tasks. The predicted properties are shape, color, and material, which are all categorical variables except for the color on Multi-dSprites. With categorical variables, the prediction accuracy is employed and compared with a baseline random guess. With the numerical one, the R2 score is shown. We report mean and stddev over 3 seeds. On Tetrominoes and Multi-dSprites, DISA correctly encodes texture and shape information into two non-overlapping subsets of its latent space dimensions. On CLEVR6 and CLEVRTex, part of the shape and texture information leaks into incorrect components.
Figure 4: Qualitative results showing the position and scale disentanglement on DISA. Given an input image (first column), we select two objects and modify their scale/position factors incrementally over 4 steps (second to fifth columns). Texture and shape are consistent across the steps, while position and scale can be independently varied correctly, suggesting that the models can confine position and scale information in their respective components. The visible limitations are displayed by ISA as well, e.g., the missing parts of an object after moving another object covering it or even the inconsistent masks produced when incrementing the scale over a certain point. Additionally, on Tetrominoes, scaling an object does not yield completely wrong results, which is not trivial as the model only sees shapes of fixed size during training.
To verify how effective DISA is in constraining the texture and shape information into the desired components, we utilize property prediction (Figure 2): we attempt to predict the ground-truth property of an object from the subset of its slot where the related information should be present, and also from a subset where the information should not be encoded. To do so, we first train multiple MLPs, each predicting one property of the objects based on the respective part of the slots, such as the shape of a tetromino given its shape vector (Figure 2(a)). In this task, we expect excellent results as we infer some property leveraging the part of the representations that should encode it. Instead, if we switched the parts and trained another set of MLPs to predict, e.g., the color of an object from its shape-related components, we would ideally expect to obtain poor results. An “inverse" property prediction task is carried out to gain insight into this aspect (Figure 2(b)). In line with previous literature (Greff et al., 2019; Dittadi et al., 2021; Papa et al., 2022), we measure performance using the accuracy for categorical variables and the coefficient of determination (R2 score) for numerical variables (i.e. only the color in Multi-dSprites). We report the mean and standard deviation of accuracy/R2 computed over the three seeds on the test set. As baselines, we provide the accuracy of a random guess for each property on every dataset. Moreover, since the objects can be represented in any order, the ground truth labels must be matched with the proper predicted slots. We do so by computing cosine similarities between the target and predicted masks and then coupling each ground truth label with the output of its most similar slot.
The results of these experiments on Tetrominoes, Multi-dSprites, CLEVR6, and CLEVRTex are reported in Figure 3. As expected, DISA reaches near-perfect property prediction scores on all datasets (except for CLEVRTex) when exploiting the correct groups of features. When inverting them, we notice on Tetrominoes and Multi-dSprites that the accuracy of the categorical predictions drops approximately to that of a random guess, while the R2 of the numerical becomes negative. The negative R2 score indicates that the inverse color predictions are worse than those of a model constantly inferring the mean target value. Therefore, these results show that, on the first two datasets, DISA is able to correctly encode texture and shape information into two non-overlapping subsets of its latent space dimensions, achieving the purpose of this work. In the case of CLEVR6, we find that although the color can be successfully restricted within the texture components, part of the shape and material information leaks respectively into the texture and shape latent features. On CLEVRTex, as both ISA and DISA are unable to precisely segment and reconstruct the objects, it is likely that the representations do not contain the necessary information to allow near-perfect accuracy even on the regular property prediction task. Even in this case, part of the shape and texture information leaks into incorrect components.
We further present in Figure 4 qualitative results showing that the position and scale disentanglement properties of DISA are preserved. Given an input image (first column), we select two objects and modify their scale or position factors incrementally over 4 steps (second to fifth columns). As visible, texture and shape are consistent across the steps, while position and scale can be independently varied with success. This suggests that the models are able to confine position and scale information in their respective components. Note that the visible limitations are displayed by ISA as well. For example, the missing parts of an object after moving another object covering it, or even the inconsistent masks produced when incrementing the scale over a certain point. Moreover, it is interesting to see that, on Tetrominoes, scaling an object does not yield completely wrong results. This is not trivial as, for a given shape, only a single variation of its scale is present in the dataset, meaning that the model only sees shapes of fixed size during training.
@@ 5.3 Compositional and Generative Capabilities
Figure 5: Compositional capabilities of DISA on (a) Tetrominoes and (b) CLEVR6. The central row contains the input image and the predicted object masks. At the top, the reconstructed image and object textures are shown (empty slots excluded), while the bottom row presents the reconstruction and object textures after interchanging texture vectors between objects. DISA shows strong compositional generalization, enabling reliable transferring of textures between objects while preserving shape, position, and scale information. Moreover, mask predictions are highly accurate with clear background separation.
@@
Figure 6: Texture generation capabilities of DISA on (a) Tetrominoes and (b) CLEVR6. The top row contains the reconstructed input image and object texture. At the bottom, the same visualization is presented after replacing the texture vectors with ones sampled from $\mathcal{N}(\mu_{\text{tex}},0.035)$, with $\mu_{\text{tex}}$ being the mean of the texture vectors encoded from the input image. DISA’s learned latent space allows us to swap in completely new textures for specific objects independently of their shape, location, and scale, demonstrating the generative capabilities of the model.
(a) Tetrominoes
(b) CLEVR6
Figure 7: Shape generation capabilities of DISA on (a) Tetrominoes and (b) CLEVR6. In each of the 2 images, the first pair of rows shows final reconstruction (middle left), object textures (first row), and shapes (second row) obtained after encoding and decoding an input image. The second pair of rows presents final reconstructions (middle left) and individual object textures (first row) and shapes (second row) after sampling new shape vectors. On Tetrominoes, the sampled shapes are roughly defined but close to those present in the training data with exceptions such as that of the cyan object on the bottom left, while on CLEVR6 the generated shapes are almost identical to those present in the training data. The difficulty of this task is mainly due to the strongly limited number of unique shapes in these synthetic datasets and to the fact that, without changing the scale of the reference frame, the generated shape is restricted to one that fits the original frame.
Finally, we also explore the compositional and generative capabilities of DISA derived from the disentanglement of texture and shape through simple qualitative experiments. Precisely, as a first investigation, for a given image from the test set we take the representations of the objects, randomly permute their texture parts, and decode the permuted representations. Ideally, the model would output a reconstruction where the shapes, positions, and sizes of the entities are preserved, while their textures are interchanged and adapted to fit the same masks. In the second experiment, we analyze whether this ability extends to the generation of new objects by providing new sampled textures or shapes to the objects in a scene. To do so, we average the texture (or shape) vectors encoded from a given image and sample from a Gaussian distribution centered on that mean vector. For additional results, including those on Multi-dSprites and CLEVRTex, please see Appendix F.2 and F.3.
In Figure 6 we show the texture generation results of DISA on Tetrominoes and CLEVR6. On Tetrominoes (Figure 6(a)), the decoded textures maintain the original shapes but switch appearance. From the first slot, it is also visible that we can generate coherent out-of-distribution textures, as the dataset only contains objects with plain colors, while the one we obtain mixes red and yellow. Concerning CLEVR6 (Figure 6(b)), we find again that DISA is able to precisely adapt the sampled textures to the fixed shapes, positions, and scales of the objects in the image. Figure 6 shows therefore that the learned representations of DISA also allow for generative capabilities: we can swap in completely new textures for specific objects independently of their shape, location, and scale.
We show in Figure 7 the shape generation results of DISA on Tetrominoes and CLEVR6. Note that this task is quite difficult: first, the number of unique shapes is strongly limited in these synthetic datasets; second, without changing the scale of the reference frame, the generated shape is restricted to one that fits the original frame. On Tetrominoes (Figure 7(a)), the sampled shapes are roughly defined but close to those present in the training data, with exceptions such as that of the cyan object on the bottom left. On CLEVR6 (Figure 7(b)) the generated shapes are almost identical to those present in the training data. A higher standard deviation during the sampling phase can lead to a larger variety in the generated shapes, while at the same time can strongly degrade their quality.
@@ Appendix A Sobel Filter
A 2-dimensional convolution operation with $\mathcal{K}_{x}$ and $\mathcal{K}_{y}$ is applied on the input image $\mathcal{X}$, producing two maps of the approximated horizontal and vertical derivatives at each pixel, respectively ${\mathcal{G}_{x}=\mathcal{K}_{x}*\mathcal{X}}$ and ${\mathcal{G}_{y}=\mathcal{K}_{y}*\mathcal{X}}$. Figure 8(b) shows an example of the resulting $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$ of the grayscale image 8(a). To obtain the final filtered input containing the detected edges, the maps are aggregated as
@@ Appendix B Datasets
Tetrominoes consists of $35\times 35$ RGB images presenting 3 “Tetris"-like blocks and a black background. The blocks are characterized by one of 6 possible colors (red, green, blue, yellow, magenta, cyan) and a shape and orientation within 19 possible. Figure 10(a) shows a few examples extracted from the dataset. The tetrominoes can appear anywhere in the image, but cannot overlap with each other. Furthermore, multiple shapes/orientations/colors of the same kind can be included in a single sample. We employ a total of 60K samples for the training set and 320 for the test set, in line with Greff et al. (2019); Locatello et al. (2020).
@@
CLEVR, introduced by Johnson et al. (2017), is a collection of $240\times 320$ images of 3D scenes containing a number of objects ranging from 3 to 10 (excluding the background). There are three types of shapes, i.e. cube, sphere, and cylinder, two possible sizes (small and large), eight colors (gray, red, blue, green, brown, purple, cyan, and yellow), and two materials (shiny “metal" and matte “rubber"). Some samples from CLEVR are presented in Figure 10(c). We also employ a filtered version of this dataset, called CLEVR6, that has a maximum of 6 entities in the scenes, background excluded. The number of training samples, in this case, is 70K, while the test samples are 15K. In all our experiments, a center crop of $192\times 192$ is applied, followed by a resize to a resolution of $128\times 128$.
@@ E.1 Training and Validation Curves
(a) Material - Training
(b) Material - Validation
@@ Appendix F Additional Experimental Results
F.1 Ablation: Impact of Variance Regularization and Sobel Filter on Texture and Shape Disentanglement
@@ F.3 Generative Results
As anticipated, we sample the new textures from a Gaussian distribution centered on the mean of the texture vectors of the objects in a scene, with a standard deviation of 0.035 (manually tuned). Similarly, we sample new shape vectors from a Gaussian distribution centered on the mean of the shape vectors of the foreground objects in a scene, with a standard deviation manually tuned. On Tetrominoes and Multi-dSprites, the standard deviation is set to 0.1, while on CLEVR and CLEVRTex 0.01.
@@
(a) Multi-dSprites
(b) CLEVRTex
Figure 37: Shape generation results on Multi-dSprites and CLEVRTex. In each of the 2 images, the first pair of rows shows final reconstruction (middle left), object textures (first row), and shapes (second row) obtained after encoding and decoding an input image. The second pair of rows presents final reconstructions (middle left) and individual object textures (first row) and shapes (second row) after sampling new shape vectors.